# ShortGPT: Layers in Large Language Models are More Redundant Than You Expect

### v2

@@ Abstract
As Large Language Models (LLMs) continue to advance in performance, their size has escalated significantly, with current LLMs containing billions or even trillions of parameters. However, in this study, we discovered that many layers of LLMs exhibit high similarity, and some layers play a negligible role in network functionality. Based on this observation, we define a metric called Block Influence (BI) to gauge the significance of each layer in LLMs. We then propose a straightforward pruning approach: layer removal, in which we directly delete the redundant layers in LLMs based on their BI scores. Experiments demonstrate that our method, which we call ShortGPT, significantly outperforms previous state-of-the-art (SOTA) methods in model pruning. Moreover, ShortGPT is orthogonal to quantization-like methods, enabling further reduction in parameters and computation. The ability to achieve better results through simple layer removal, as opposed to more complex pruning techniques, suggests a high degree of redundancy in the model architecture.
Figure 1: The results on MMLU by pruning 25% parameters with different methods.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains. Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research has tended to increase model parameters consistently to boost performance. However, the resulting massive models containing billions or even trillions of parameters pose stringent hardware requirements, hindering their practical deployment and use. To alleviate the demanding hardware requirements for deploying massive trained models, many researchers have focused on model compression techniques (Zhu et al., 2023) that reduce the inference cost. Model compression methods can be broadly categorized into two categories: quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022, 2024) and pruning (LeCun et al., 1989; Han et al., 2015). Quantization methods quantize the weights and activations to lower precision. However, the acceleration benefits of quantization are dependent on hardware support and sometimes require additional fine-tuning to maintain good performance. In contrast, pruning methods remove redundant model parameters to reduce the overall parameter count. Pruning can be directly applied to a trained model without retraining and is generally more hardware-friendly than quantization approaches. While recently there has been significant progress in large language model compression methods based on model pruning, existing methods are often designed to be relatively complex. Some require the use of gradient information (Zhang et al., 2023) or only focus on compression in width (Ashkboos et al., 2024), rendering them overly sophisticated for practical applications. Therefore, there is a need to explore simple and efficient model pruning approaches tailored specifically for large language models. Identifying effective model pruning techniques necessitates studying model redundancy (Huang et al., 2021; Dalvi et al., 2020). Previous research on model redundancy has typically focused on relatively smaller models, such as convolutional neural networks (CNNs) or small-sized Transformers. For model pruning in LLMs, most previous works have concentrated on tensor-wise redundancy analysis, investigating the redundancy within each parameter tensor. However, a key finding in this paper is that LLMs exhibit significant redundancy at the layer level, enabling the simple removal of entire layers without substantially impacting downstream task performance. For instance, when deleting the last 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model, the results on the MMLU benchmark (Hendrycks et al., 2020) only drop from 55.0 to 52.2 (a 95% retention). Furthermore, by removing the last 22 layers (55% of the total 40 layers), resulting in a 5.6 billion parameter model, we can still obtain a score of 47.2 on MMLU without any fine-tuning, even outperforming the LLaMA 2-7B model. In this paper, we propose to analyze the layer-wise redundancy through the lens of Block Influence (BI), which measures the hidden states transformations during the modeling process of LLMs. We find BI is a more relevant indicator of the layer’s importance in LLMs, and we can perform model pruning by simply deleting redundant layers with BI. This straightforward act of removing specific layers significantly outperforms previous more complex pruning methods. Our findings highlight substantial redundancy in current LLM architectures and shed light on opportunities for more efficient LLM training in the future. The main contributions of our paper are summarized as follows:
@@
  • We analyze the redundancy in large language models (LLMs) and find that they exhibit significant redundancy at the layer level. This insight inspires us to prune LLMs by simply removing redundant layers.
  • We propose a metric called Block Influence (BI) as an effective indicator of layer importance. Through a quantitative analysis, we demonstrate that LLMs possess redundancy both in depth (layers) and width (parameters within layers).
@@
  • Based on the BI metric, we propose a simple yet effective pruning strategy by removing layers with low BI scores. Experimental results show that our method maintains 92% performance while reducing approximately 25% of the parameters and computation, outperforming previous state-of-the-art methods.
@@
(a) Layer pruning order. We first calculate the BI score of each layer in LLM, and then prune the redundant layer in ascending BI score.
(b) Performance of ShortGPT with increasing pruning ratio.
Figure 2: Our Method: firstly, we sort layers by their BI score (introduced in Sec. LABEL:layerimportacne), and then delete layers in ascending order. All results are based on Llama2-7B-Base.
@@ 2 Methodology
In this section, we present the methodological framework of our layer deletion approach for LLMs, elucidating the underlying principles and techniques employed. We first quantify the layer redundancy problem present in current prominent LLMs, such as LLaMA 2 and Baichuan 2. Then, we introduce a metric called Block Influence (BI), designed to assess the transformation of hidden states by each layer during LLM inference. Building upon BI, we apply layer deletion to LLMs, thereby reducing their inference cost without compromising their predictive accuracy or linguistic capabilities.
2.1 Layer redundancy
The predominant LLMs in current are primarily based on the Transformer (Vaswani et al., 2017). The transformer architecture is based on an attention mechanism, typically consisting of several residual layers stacked upon each other. Transformer is a sequence to sequence mapping, which can be defined as $y=f(X,\theta)$, where $X\in R^{T\times n}$, $Y\in R^{T\times n}$, $T$ is the length of the sequence, $n$ is the vocabulary size, $\theta$ is the learnable parameters. The formal expression of an L-layer transformer is as follows:
$\textbf{\footnotesize{X}}_{0}$ $=\textbf{\footnotesize{X}}\textbf{W}_{\text{emb}}$
$\textbf{\footnotesize{A}}_{i+1}$ $=\text{ATTN}(\text{LN}(\textbf{\footnotesize{X}}_{i}))+\textbf{\footnotesize{X}}_{i}$
$\textbf{\footnotesize{X}}_{i+1}$ $=\text{FFN}(\text{LN}(\textbf{\footnotesize{A}}_{i+1}))+\textbf{\footnotesize{A}}_{i+1}$
Y $=\textbf{\footnotesize{X}}_{L}\textbf{W}_{\text{head}}$ (1)
where $\textbf{W}_{\text{emb}}\in R^{n\times d}$ is the word embedding matrix, $\textbf{W}_{\text{head}}\in R^{d\times n}$ is the output projection matrix of the transformer, which are sometimes tied with the $\textbf{W}_{\text{emb}}$ (Chowdhery et al., 2023), $d$ is the hidden dim of the transformer. ATTN refers to the attention layer and FFN means the feed-forward layers, $\textbf{X}_{i}\in R^{T\times d}$ is the hidden states of the $i^{th}$ layers. Given that the transformer structure is composed of several identical layers, a natural question is what the differences and connections between the functionalities and roles of these identical layers are. Previous works found that the Transformer possesses certain semantic capabilities in earlier layers (Hasan et al., 2021). In this work, we uncover a significant level of redundancy between the layers of the transformer.
Layer Redundancy: A network with high redundancy should contain some redundant layers, which have a minimal impact on the network’s final performance. This may be due to these layers having homogenized functionalities compared to other layers in the network. We have found a high degree of redundancy in the current LLMs by omitting specific layers.
(a) Llama2-7B-Base
(b) Baichuan2-7B-Base
Figure 3: Performance of removing single layer.
We observed that in many cases, the network’s prediction of $Y$ remains largely unchanged. This phenomenon demonstrates a significant level of redundancy across the layers of the transformer. Figure 3 shows the perplexity and MMLU (Hendrycks et al., 2020) score of omitting a certain layer in the Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese. For more benchmark results, please refer to the section 3. Figure 3 reveals that some layers do not play a crucial role in LLMs. Dropping a certain layer from the LLMs may have a minimal effect on the final results, and the same holds for the perplexity. Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the final layer often being more critical.
Figure 4: Illustration of our Block Influence(BI) metric. We omit LayerNorm for visual conciseness.
@@ 2.2 Layer importance
As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers. Notably, the deeper layers demonstrate a higher level of redundancy. To remove these redundant layers, we need a model-intrinsic metric to measure the importance of the layer. An intuitive method is to use the magnitude of each layer’s output as a measure of its importance, as a larger magnitude implies more activated neurons. In (Samragh et al., 2023), they proposed to use relative magnitude $||\frac{f(x)}{x+f(x)}||$ to measure the importance of layers. To characterize the influence of a certain layer. We defined a new metric, Block Influence (BI), under the assumption that the more a transformer block changes the hidden states, the more influential this layer is. As Figure 4 shows, the BI score of $i^{th}$ blocks can be calculated as follows:
@@
where $X_{i,t}$ means the $t^{th}$ row of $X_{i}$. Our empirical evidence supports that BI effectively reflects the importance of a layer. Figure 8 shows these different metrics in detail.
@@ 2.3 Layer Removal
The method we propose is straightforward: layer removal, which we simply referred as ShortGPT. We delete certain layers in LLMs based on BI score. Figure 2 illustrates our pruning method. First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 (Rae et al., 2019). Then we collect the hidden states of each layer during inference on these samples. Next, we calculate the BI score based on the collected hidden states. Finally, we sort layers in ascending order according to the BI, and delete the layers with the smaller importance. The number of layers to be deleted can vary to trade off the speed and performance. The details of our layer remove setting can be found in Appendix B
@@ 3.1.1 Models
3.1.2 Benchmarks
In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted evaluations on the most commonly used Benchmark MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024) for evaluating large models. In addition, we also followed LaCo (Yang et al., 2024) to evaluate a wider dataset.
MMLU (Hendrycks et al., 2020) is a benchmark aimed at measuring the knowledge acquired during pre-training by specifically evaluating models in zero-shot and few-shot settings. This makes benchmarks more challenging and similar to the way we evaluate humans. This benchmark covers 57 subjects including STEM, humanities, social sciences, etc. Its difficulty ranges from beginner to advanced professional level, and it tests world knowledge and problem-solving ability.
CMMLU (Li et al., 2024) is a comprehensive Chinese language assessment dataset designed specifically to evaluate LLM’s advanced knowledge and reasoning abilities in the context of Chinese language and culture. CMMLU covers 67 topics, from elementary school to university or professional level. Including natural sciences, as well as humanities and social sciences, it also includes many contents with Chinese characteristics.
CMNLI (Xu et al., 2020) is part of the Chinese language understanding assessment benchmark. It consists of two parts: XNLI and MNLI. HellaSwag (HeSw) (Zellers et al., 2019) is a challenging dataset for evaluating commonsense NLI that is especially hard for state-of-the-art models, though its questions are trivial for humans. PIQA (Bisk et al., 2020) is a multi-choice question and answer dataset that focuses on daily scenarios. This dataset explores the model’s grasp of the laws of the real physical world through daily scenarios. CHID (Zheng et al., 2019) is an idiom cloze test dataset that mainly focuses on the selection of candidate words and the representation of idioms. CoQA (Reddy et al., 2019) is a large-scale dataset used for conversational question-answering tasks, containing over 127000 questions and their corresponding answers. BoolQ (Clark et al., 2019) is a question-answer dataset containing 15942 examples of yes/no questions. These problems occur naturally - they are generated in an environment that is silent and unconstrained. Race (Lai et al., 2017) is a large-scale reading comprehension dataset collected from English examinations in China, which are designed for middle school and high school students. XSum(Hasan et al., 2021) is used to evaluate abstract single document summarization systems. The goal is to create a short, one-sentence new summary of what the article is about. C3 (Sun et al., 2020) is a machine reading comprehension dataset with multiple choices, consisting of multiple-choice questions, reading materials from Chinese proficiency exams, and ethnic Chinese exams. PG19 (Rae et al., 2019) is a long document dataset from books used to test the effectiveness of language modeling.
3.1.3 Baselines
@@
To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models followed by LaCo (Yang et al., 2024). For our method, we use PG19 for layer importance and perplexity calculation.
@@
LLMPru. (Ma et al., 2024) adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. LLMPru. applies post training to the pruned model, but for fair comparison, we do not apply post training to it.
@@
SliceGPT (Ashkboos et al., 2024) is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network. Specifically, they applied PCA to the hidden representation from shallow to deep layers, and incorporated the dimension reduction matrix into existing network parameters.
@@
LaCo (Yang et al., 2024) is a pruning method for large language models based on reducing layers. LaCo gradually merges similar layers from deep to shallow and sets a threshold to avoid continuously merging too many layers.
@@
Llama2-7B Dense 0.00% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92 44.52 100.00
LLMPrun. 27.0% 34.33 56.46 71.22 25.25 36.54 42.51 55.20 22.56 22.35 11.51 25.64 23.33 25.25 32.84 73.76
SliceGPT 26.4% 31.70 50.27 66.21 20.79 36.54 41.36 38.32 21.07 21.66 4.89 39.78 28.92 25.37 32.84 73.76
LaCo 27.1% 34.43 55.69 69.80 36.14 40.38 45.70 64.07 22.61 23.61 15.64 39.67 26.45 25.24 38.41 86.28
ShortGPT 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.71 32.25 35.17 0.67 39.62 43.96 32.25 42.60 95.69
Llama2-13B Dense 0.00% 32.99 74.78 79.71 47.35 50.00 66.91 82.39 57.95 60.38 23.45 47.51 55.00 38.40 51.91 100.00
LLMPrun. 24.4% 33.03 67.76 76.66 35.64 40.38 50.86 56.42 22.47 22.08 19.17 32.33 25.21 24.71 38.97 75.07
SliceGPT 23.6% 29.82 55.71 69.04 19.31 36.54 47.26 37.86 23.41 24.03 5.27 41.92 37.14 25.79 34.84 67.11
LaCo 24.6% 32.86 64.39 74.27 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 48.30 93.05
ShortGPT 24.6% 33.00 66.64 73.45 36.61 50.00 58.64 62.48 58.35 60.17 17.59 46.90 54.69 38.38 50.53 97.34
Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 26.96 24.09 20.82 64.55 53.87 56.95 53.63 100.00
LLMPrun. 24.2% 32.28 53.66 71.82 69.80 53.85 47.83 61.19 21.96 22.28 15.98 41.64 24.93 25.69 41.76 77.87
SliceGPT 22.2% 32.07 25.29 50.33 14.85 36.54 19.57 39.30 23.53 22.49 0.00 26.58 25.18 25.25 26.23 56.38
LaCo 24.2% 33.00 52.28 68.50 76.24 42.31 47.26 56.15 28.99 27.72 12.03 50.85 31.53 31.24 42.93 80.05
ShortGPT 24.2% 33.30 56.96 67.68 65.63 50.00 46.70 67.83 53.26 46.76 0.04 56.33 45.77 47.87 49.08 91.52
@@
ShortGPT 24.7% 32.81 60.55 71.60 80.17 47.13 54.30 62.54 55.77 56.41 15.14 60.16 52.11 58.86 54.43 87.33
Table 1: Comparison of pruning methods on multiple natural language benchmarks. The results of LLMPrun., SliceGPT and LaCo are reported from LaCo.

### v3

@@ Abstract
As Large Language Models (LLMs) continue to advance in performance, their size has increased significantly, with current LLMs containing billions or even trillions of parameters. In this study, we identify notable redundancy across the layers of LLMs, where some layers contribute minimally to overall network functionality. To quantify this, we introduce a metric called Block Influence (BI) which use the similarity between layer’s input and output to measure the importance of each layer. Based on the observation of layer redundancy, we propose a straightforward pruning method: layer removal, which eliminates redundant layers based on their BI scores. Our approach, termed ShortGPT, demonstrates superior performance over previous state-of-the-art pruning methods. Moreover, ShortGPT is orthogonal to quantization-like methods, enabling further reduction in parameters and computation. The ability to achieve better results through simple layer removal, as opposed to more complex pruning techniques, suggests a high degree of redundancy across layers, not only in transformer models but also in non-transformer models. We hope this work will contribute to future research in LLM compression.
@@ 1 Introduction
The field of large language models (LLMs) has witnessed rapid development recently, with LLMs achieving impressive performance across various domains. Guided by the scaling laws identified in prior work (Kaplan et al., 2020; Hoffmann et al., 2022), current LLM research tend to increase model parameters to boost performance. As a result, modern LLMs, which can comprise billions to trillions of parameters, require significant hardware resources for deployment, creating substantial barriers to their practical use.
To mitigate the hardware demands of large models, model compression techniques have become a critical area of focus (Zhu et al., 2023). These techniques are generally divided into quantization (Liu et al., 2021; Gholami et al., 2022; Dettmers et al., 2022; 2024) and pruning(LeCun et al., 1989; Han et al., 2015; Frantar & Alistarh, 2023). Quantization reduces the precision of model parameters, but its effectiveness often requires specific hardware support. In contrast, pruning method removes redundant parameters to decrease the model’s size and computation, offering a more flexible and hardware-agnostic approach. Despite its advantages, many existing pruning methods are complex; for example, some require gradient information (Ma et al., 2024), which limits their practicality.
In this paper, we focus on the issue of layer redundancy in LLMs and propose a novel approach for simplifying these models. We introduce Block Influence (BI), a metric that quantifies how much the hidden state changes after passing through each layer, providing a more direct measure of a layer’s importance. Leveraging this insight, we propose a simple yet effective pruning method ShortGPT, which identifies and removes layers with lower BI scores, significantly reducing model size without sacrificing much performance.
To evaluate our approach, we conducted evaluation across comprehensive benchmarks. Our experiments revealed that our method exhibits a smaller performance decrement compared to the previous methods. For instance, removing 10 layers (25% of the total 40 layers) from the LLaMA 2-13B model resulted in only a slight drop in performance on the MMLU benchmark (Hendrycks et al., 2020), from 55.0 to 52.2. Our findings highlight substantial redundancy in current LLMs and suggest potential avenues for improving the efficiency of model training by reducing inherent redundancy in the future.
The main contributions of our paper are summarized as follows:
@@
  • We analyze the redundancy in large language models (LLMs) and find that they exhibit significant redundancy at the layer level. This finding inspire us to prune LLMs by simply removing redundant layers.
@@
  • We propose a metric called Block Influence (BI) as an indicator of layer importance. Based on BI, our layer removal method maintains approximately 90% performance while reducing approximately 25% of parameters, outperforming previous state-of-the-art methods.
@@
(a) Perplxity
(b) MMLU
Figure 1: Performance of removing certain layer from LLMs. We can see that certain layers are redundant, and their removal results in minimal performance degradation.
2 Motivation
2.1 Background
The predominant LLMs are primarily based on the Transformer architecture (Vaswani et al., 2017), with the pre-norm configuration being the most commonly adopted, as in models like LLaMA (Touvron et al., 2023). The pre-norm configuration, where layer normalization is applied before the self-attention and feed-forward layers, offers several advantages such as faster convergence, improved training stability, and better scalability for deeper networks (Xiong et al., 2020; Liu et al., 2020; Wang et al., 2024). Due to these benefits, the pre-norm approach has been adopted even in non-transformer models, such as Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023). For the sake of simplicity in descriptions, our analysis primarily focuses on the Transformer architecture, though we extend our experiments to non-Transformer structures in Section 4.4.
Figure 2: The cosine similarity between a layer’s input and output during the training process. The horizontal axis (X-axis) represents the number of training tokens, while the vertical axis (Y-axis) depicts the degree of similarity. Notably, the model employing post-normalization exhibits divergence after approximately $\sim$26B tokens of training. Training setting is provided in E.
However, we observe that when pre-norm is adopted, the similarity between the input and output of transformer layers tends to be higher, as illustrated in Figure 2. This high similarity indicates that certain layers induce minimal changes to the hidden states, suggesting they contribute little to the model’s overall function. A detailed mathematical explanation for this phenomenon is provided in Appendix A. Which suggests that the deep layers of the model with pre-norm might not play a critical role in the overall function, and that the layers in large language models could be more redundant than expected, which motivates the layer-removal based pruning method we explore in the next section.
2.2 Layer redundancy
Table 1: Ablation of removing FFN and Attention of Llama2-7B-Base. We sample 100 instances from PG19 (Rae et al., 2019) to calculate PPL.
Delete PPL
None 7.60
The whole last layer 13.37
Attention of the last layer 7.65
FFN of the last layer 12.35
As discussed in the previous section, we speculate that the LLMs exhibit layer redundancy. To verify this, we assess the performance degradation caused by removing individual layers of two popular models, Llama2-7B-Base (Touvron et al., 2023), an English based LLMs, and Baichuan2-7B-Base (Yang et al., 2023) which is mainly focused on Chinese. Figure 1 confirms our speculation, which reveals that some layers do not play a crucial role in LLMs, causing little degradation when omitting them individually. Moreover, this redundancy is primarily manifested in the middle to later layers of the network, with the initial layers and the last layer often being more critical. Notably, we found the last layer to be particularly important, aligning with findings from LLM Pruner (Ma et al., 2024). This observation contradicts our mathematical explanation in Appendix A which suggests that deeper layers tend to be more redundant. We posit that this discrepancy arises because the final FFN effectively functions as part of the token classifier and should be considered in conjunction with the language model head.To verify our hypothesis, we conducted further investigation, detailed in Table 1. The results show that within the last layer, the FFN component is crucial, while the Attention module is less significant. This finding supports our interpretation of the final layer’s importance.
@@ 3 Methodology
In this section, we present the methodological framework of our layer removal approach for LLMs, elucidating the underlying principles and techniques employed. We begin by introducing Block Influence (BI), a novel metric designed to assess the hidden states transformation of each layer. Leveraging BI, we then detail our layer removal method.
@@ 3.1 Layer importance
As outlined in the preceding section, the layers of LLMs exhibit redundancy, with varying degrees of redundancy across different layers. To capture this, we introduce a new metric, Block Influence (BI), to measure the degree of transformation performed by each layer. The BI score of $i^{th}$ layer can be calculated as follows:
@@
where $X_{i,t}$ means the $t^{th}$ row of hidden states of $i^{th}$ layer. Lower BI score imply that $X_{i}$ and $X_{i+1}$ exhibit high cosine similarity, suggesting that the layer makes minimal transformations to the hidden states and is therefore less important. We plot the BI scores of a single layer and the PPL after removing it separately, as shown in the Figure 3. The results demonstrate a positive correlation between the BI score and the importance of a layer.
(a) Llama2 7B
(b) Baichuan2 7B
Figure 3: The BI score of a layer and the PPL after removing the layer.
@@ 3.2 Layer Removal
Our goal is to obtain a pruned model that remains as close as possible to the original model. Since an LLM functions as a series of transformations applied to hidden states across its layers and we can determine the importance of each layer, we propose a straightforward pruning method: layer removal, which we refer to as ShortGPT. We delete certain layers in LLMs based on BI score. First of all, we construct a calibration set, which is a set of unlabelled text samples such as PG19 (Rae et al., 2019). Then we collect the hidden states of each layer during inference on these samples. Next, we calculate the BI score based on the collected hidden states. Finally, we sort layers in ascending order according to the BI, and delete the layers with the lower BI score. The number of layers to be deleted can vary to trade off the speed and performance. The details of our layer removal setting can be found in Appendix D.
@@ 4.1 Experimental Setup
Benchmarks.
In order to comprehensively evaluate the changes in the ability of large language models before and after pruning, we conducted comprehensive evaluation from five aspect: Reasoning: CMNLI (Li et al., 2024), HellaSwag (HeSw) (Zellers et al., 2019), PIQA (Bisk et al., 2020). Language: CHID (Zheng et al., 2019), WSC (Levesque et al., 2012). Knowledge: CommonSenseQA (CoQA) (Reddy et al., 2019), BoolQ (Clark et al., 2019). Examination: MMLU (Hendrycks et al., 2020), CMMLU (Li et al., 2024). Understanding: Race-High/Middle (H/M) (Lai et al., 2017), XSum (Hasan et al., 2021), C3 (Sun et al., 2020) and PG19 (Rae et al., 2019). For more details, please refer to Appendix G
Baselines.
@@
To evaluate the effectiveness of our method, we compared several structured pruning methods for large language models, including:
@@
1) LLMPru (Ma et al., 2024), which adopts structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the majority of the LLM’s functionality. LLMPru. applies post training to the pruned model, but for fair comparison, we do not apply post training to it.
@@
2) SliceGPT (Ashkboos et al., 2024), which is a post-training sparsification scheme that replaces each weight matrix with a smaller matrix, reducing the embedding dimension of the network. Specifically, they applied PCA to the hidden representation from shallow to deep layers, and incorporated the dimension reduction matrix into existing network parameters.
@@
3) LaCo (Yang et al., 2024), which is a pruning method for large language models based on reducing layers. LaCo gradually merges similar layers from deep to shallow and sets a threshold to avoid continuously merging too many layers.
For our evaluation, we use PG19 for layer importance and perplexity calculation. The models, baselines and evaluate benchmarks is the same as LaCo.
Table 2: Comparison of pruning methods on multiple natural language benchmarks. The results of LLMPrun., SliceGPT and LaCo are reported from LaCo. The last column reports the relative performance retention.
@@
Llama2-7B Dense 0.00% 32.99 71.26 77.91 41.66 50.00 64.62 71.62 35.71 34.19 19.40 43.56 45.39 32.92 47.78 100.00
LLMPrun. 27.0% 34.33 56.46 71.22 25.25 36.54 42.51 55.20 22.56 22.35 11.51 25.64 23.33 25.25 34.78 72.79
SliceGPT 26.4% 31.70 50.27 66.21 20.79 36.54 41.36 38.32 21.07 21.66 4.89 39.78 28.92 25.37 32.84 68.73
LaCo 27.1% 34.43 55.69 69.80 36.14 40.38 45.70 64.07 22.61 23.61 15.64 39.67 26.45 25.24 38.41 80.39
ShortGPT 27.1% 32.95 53.02 66.43 24.68 52.46 47.99 74.71 32.25 35.17 0.67 39.62 43.96 32.25 41.24 86.31
Llama2-13B Dense 0.00% 32.99 74.78 79.71 47.35 50.00 66.91 82.39 57.95 60.38 23.45 47.51 55.00 38.40 55.14 100.00
LLMPrun. 24.4% 33.03 67.76 76.66 35.64 40.38 50.86 56.42 22.47 22.08 19.17 32.33 25.21 24.71 38.97 70.67
SliceGPT 23.6% 29.82 55.71 69.04 19.31 36.54 47.26 37.86 23.41 24.03 5.27 41.92 37.14 25.79 34.85 63.20
LaCo 24.6% 32.86 64.39 63.20 40.10 52.88 52.66 63.98 54.49 56.55 14.45 44.93 45.93 32.62 47.62 86.36
ShortGPT 24.6% 33.00 66.64 73.45 36.61 50.00 58.64 62.48 58.35 60.17 17.59 46.90 54.69 38.38 50.53 91.64
Baichuan2-7B Dense 0.00% 33.37 67.56 76.17 85.56 50.00 63.14 74.10 52.63 51.04 20.82 64.55 53.87 56.95 57.67 100.00
LLMPrun. 24.2% 32.28 53.66 71.82 69.80 53.85 47.83 61.19 21.96 22.28 15.98 41.64 24.93 25.69 41.76 72.41
SliceGPT 22.2% 32.07 25.29 50.33 14.85 36.54 19.57 39.30 23.53 22.49 0.00 26.58 25.18 25.25 26.23 45.48
LaCo 24.2% 33.00 52.28 68.50 76.24 42.31 47.26 56.15 28.99 27.72 12.03 50.85 31.53 31.24 42.93 74.44
ShortGPT 24.2% 33.30 56.96 67.68 65.63 50.00 46.70 67.83 53.26 46.76 0.04 56.33 45.77 47.87 49.08 85.10
@@
ShortGPT 24.7% 32.81 60.55 71.60 80.17 47.13 54.30 62.54 55.77 56.41 15.14 60.16 52.11 58.86 54.43 87.35